QUESTION 3 The motion of an object where the resultant force acting on it is zero, is described by Newtons 3rd Law Newtons
Answer #1

Remember that Newton's 1st law says that any object will remain at rest or continue moving, until external force is applied.

Newton's 2nd law: F_net = m*a

So here given scenario cannot be described by Newton's 1st law, since external force is applied, But since also given that net force on object is zero, So

F_net = 0

m*a = 0

a = 0

a = acceleration = dV/dt = 0

So, V = Constant

So object will keep moving with constant speed

So above situation can be explained by Newton's 2nd law.

Correct option is D.

Let me know if you've any query.